OpenStudy (anonymous):

the geometric mean of two numbers is 2\sqrt{5} one of the numbers is 6. find the other number *******metal ************

OpenStudy (johnweldon1993):

Okay...well to find the geometric mean...you take the two numbers.....place them under a square root sign (since it is 2 numbers)...and multiply them.. so \[\sqrt{6 \times x} = \frac{ 2 }{ \sqrt{5} }\] So like the last time....lets square both sides to get rid of the square root signs \[\sqrt{6 \times x}^2 = (\frac{ 2 }{ \sqrt{5} })^2\] becomes \[6x = \frac{ 4 }{ 5 }\] Lets multiply both sides by 5 \[30x = 4\] And divide both sides by 30 to solve for 'x' \[x = \frac{ 4 }{ 30 }\] and simplify \[x = \frac{ 2 }{ 15 }\] Now lets check by putting it in the original equation \[\sqrt{6 \times \frac{ 2 }{ 15 }} = \frac{ 2 }{ \sqrt{5} }\] \[\sqrt{\frac{ 12 }{ 15 }} = \frac{ 2 }{ \sqrt{5} }\] this can be written as \[\frac{ \sqrt{4}\sqrt{3} }{ \sqrt{5}\sqrt{3} } = \frac{ 2\sqrt{3} }{ \sqrt{5}\sqrt{3} } = \frac{ 2 }{ \sqrt{5} }\] so yes it is correct....the other number is 2/15